Find the area of a regular octagon with a 18.7 m apothem and a side length of 15.5 m
Which answer is it?
a. 145.0 m^2
b. 1059.4 m^2
c. 1159.4 m^2
d. 2318.8 m^2

Answer by jim_thompson5910(35256)

Imagine you cut the octagon into triangles by drawing segments from the center to the vertices.

Area of one triangle (in the octagon):

A = (bh)/2

A = (15.5*18.7)/2

A = 144.925

So one triangle has an area of 144.925 m^2. This means that 8 of these equal triangles have a total area of 8*144.925 = 1159.4 m^2